OpenStudy (anonymous):

Solve the following system of equations. 4x - 2y - z = -5 x - 3y + 2z = 3 3x + y - 2z = -5 a (0, 1, 3) b (0, -1, 3) c (0, 1, -3) d (0, -1, -3)

OpenStudy (amoodarya):

you can find x from 2nd equation and put it into others now now you have 2 variable to find so can do it again to find y or z then put that variable to find other variable , then like that find x

OpenStudy (amoodarya):

4x - 2y - z = -5 x - 3y + 2z = 3 ---> x=3y-2z+3 3x + y - 2z = -5 4(3y-2z+3)-2y-z=-5 3(3y-2z+3)+y-2z=-5

OpenStudy (anonymous):

so -5 would be x?

OpenStudy (anonymous):

Sorry its been w while since iv done this.

OpenStudy (amoodarya):

4(3y-2z+3)-2y-z=-5 12y-8z+12-2y-z=-5 3(3y-2z+3)+y-2z=-5 9y-6z+9+y-2z=-5 10y-9z=-5-12 10y-5z=-5-9

OpenStudy (amoodarya):

if you try x=0 y=1 z=3

OpenStudy (anonymous):

sorry OS is making my computer freez

OpenStudy (anonymous):

freeze*